OpenStudy (megannicole51):

determine if this sequence converges absolutely, conditionally or diverge? (1)^n * n/n^1+1 i know its an alternating series

OpenStudy (megannicole51):

@hartnn

OpenStudy (agent0smith):

i can't tell what it is...?\[\large \frac{ (-1)^n n }{ n^2+1 }\]

OpenStudy (anonymous):

seems unlikely it says \(\frac{(1)^nn}{n^1+1}\)

OpenStudy (agent0smith):

I just assumed the 1 exponent was a typo

OpenStudy (anonymous):

i meant it seems unlikely that that is what the question is maybe \[\frac{(-1)^nn}{n+1}\] or what you wrote \[\frac{(-1)^nn}{n^2+1}\]

OpenStudy (megannicole51):

its n^2 it was a typo...brb

OpenStudy (anonymous):

as it alternates, all you have to check is that \[\lim_{n\to \infty}\frac{n}{n^2+1}=0\] which it is

OpenStudy (megannicole51):

simplified to 1/n which diverges

OpenStudy (agent0smith):

No but it's the terms that converge to zero (1/n) AND the fact it's alternating, that makes it convergent. Without the (-1)^n it would be divergent.

OpenStudy (agent0smith):

Think of it like this... (basic example) 0.9+0.8+0.7+0.6+0.5+0.4+0.3+0.2+0.1 = 4.5 but 0.9-0.8+0.7-0.6+0.5-0.4+0.3-0.2+0.1 = 0.5

OpenStudy (agent0smith):

the alternating adding/subtracting is what makes it converge.

OpenStudy (agent0smith):

eg 1/n vs (-1)^n/n: 1/n diverges http://www.wolframalpha.com/input/?i=sum+of+1%2Fn+ (-1)^n/n converges http://www.wolframalpha.com/input/?i=sum+of+%28-1%29%5En%2Fn+